OpenStudy (anonymous):

2a/3+1/5=7/6 help please!! solve for A!

OpenStudy (anonymous):

1.) 2a/3+1/5-7/6=0 subtract 7/6 to make the equation =0 2.) Get like denominators 3*10,5*6,6*5 have a multiple of 30 20a/30+6/30-35/30=0 3.) perform order of operations 6/30-35/30=-29/30 4.) 20a/30-29/30=0 multiply 30 to cancel denominator 30( 2a+-29)=0*30 5.) Divide 20a to -29 20a/20a-29/20a so answer is -29/20a